OpenStudy (anonymous):

find the difference quotient of f that is find f(x+h)-f(x)/h for f(x) =x^2-2x+4

OpenStudy (anonymous):

i think, it's not f(x+h)=f(x)/h but, f(x+h)-f(x)/h

OpenStudy (anonymous):

yes im sorry i wrote that wrong it is f(x+h)-f(x)/h

OpenStudy (noelgreco):

What don't you get?

OpenStudy (campbell_st):

ok will substitute x + h for x thats what f(x + h) is asking you to do \[\frac{(x + h)^2 -2(x +h) + 4 - (x^2 -2x + 4)}{h}\] the next step is to expand and then collect like terms... this will leave the numerator with terms involving h which will be able to cancel with the denominator